OpenStudy (anonymous):

Solve for X: 2/5 + 3/5x = +5/10

OpenStudy (anonymous):

in the second term, is x in the numerator or denominator? if its in the numerator then its quite simple. you get (5/5)x=5/10. therefore x=1/2

OpenStudy (anonymous):

Assuming you mean \[{2 \over 5} + {3 \over 5}x = {5 \over 10}\] I would multiply the entire equation by 5 just to keep things easier to solve

OpenStudy (anonymous):